OpenStudy (anonymous):

let g(x)=x^2-3x+8.find the value of 2[g(2)] 6 4 18 12

jimthompson5910 (jim_thompson5910):

first you need to find g(2)

jimthompson5910 (jim_thompson5910):

you do so by replacing every copy of x with 2 and evaluating g(x)=x^2-3x+8 g(2)=(2)^2-3*2+8 g(2) = ???

OpenStudy (anonymous):

18

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

g(2)=(2)^2-3*2+8 g(2)=4-6+8 g(2) = ??

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

now multiply both sides by 2

jimthompson5910 (jim_thompson5910):

g(2) = 6 2*g(2) = 2*6 2*g(2) = 12

OpenStudy (anonymous):

that equals 12 so the answer is 12?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

okay thanks

jimthompson5910 (jim_thompson5910):

yw